### v2

# Modelling Multimodal Integration in Human Concept Processing with Vision-Language Models

Abstract
Text representations from language models have proven remarkably predictive of human neural activity involved in language processing, with the recent transformer-based models outperforming previous architectures in downstream tasks and prediction of brain responses. However, the word representations learnt by language-only models may be limited in that they lack sensory information from other modalities, which several cognitive and neuroscience studies showed to be reflected in human meaning representations. Here, we leverage current pre-trained vision-language models (VLMs) to investigate whether the integration of visuo-linguistic information they operate leads to representations that are more aligned with human brain activity than those obtained by models trained with language-only input. We focus on fMRI responses recorded while participants read concept words in the context of either a full sentence or a picture. Our results reveal that VLM representations correlate more strongly than those by language-only models with activations in brain areas functionally related to language processing. Additionally, we find that transformer-based vision-language encoders—e.g., LXMERT and VisualBERT—yield more brain-aligned representations than generative VLMs, whose autoregressive abilities do not seem to provide an advantage when modelling single words. Finally, our ablation analyses suggest that the high brain alignment achieved by some of the VLMs we evaluate results from semantic information acquired specifically during multimodal pretraining as opposed to being already encoded in their unimodal modules. Altogether, our findings indicate an advantage of multimodal models in predicting human brain activations, which reveals that modelling language and vision integration has the potential to capture the multimodal nature of human concept representations.
footnotetext: *   Equal contributionfootnotetext: †  Corresponding author
1 Introduction
How does the brain represent semantic knowledge? Whereas many computational models of language still build on the idea that meaning can be extracted from text corpora (Piantadosi and Hill, 2022), increasing evidence from cognitive science and neuroscience suggests that human semantic representations are in fact grounded in sensory experiences. Theories of embodied simulation argue that comprehending and producing words involves both sensory-motor and linguistic information (Louwerse, 2011; Barsalou, 1999; Harnad, 1990; Bergen, 2012). This view is supported by developmental studies on word meaning acquisition (Landau, Smith, and Jones, 1998; Bornstein et al., 2004) as well as behavioural and neuroimaging experiments highlighting the role of perceptual attributes in processing word meanings (Barsalou, 2008). The need to ground meaning representations in the physical world has also been emphasized by researchers in natural language processing as an important step towards better language understanding in machines (Bender and Koller, 2020; Bisk et al., 2020).
In human concept processing, visual and linguistic stimuli have been shown to mutually influence each other (Boutonnet and Lupyan, 2015; Noorman, Neville, and Simanova, 2018; Huettig, Rommers, and Meyer, 2011) and also evoke shared neural activity patterns, thought to reflect some form of abstracted, modality-independent conceptual representations (Devereux et al., 2013; Simanova et al., 2014; Popham et al., 2021; Kaup et al., 2024). Although the existence of such modality-independent activity patterns does not necessarily tell us much about their representational content, recent evidence indicates that they indeed encode both linguistic and visual information (Dirani and Pylkkänen, 2024). For example, Tang et al. (2024) showed that brain activity in visual processing areas can be successfully predicted by computational models trained to predict the same kind of responses during language processing; the effect is even stronger when such models integrate stimulus features from both modalities. Such analyses, comparing between brain activity patterns and representations generated by deep neural network (DNN) models, offer a promising way to shed further light on the nature of multimodal integration in the brain.
In this paper, we investigate the ability of vision-language models (VLMs) to capture multimodal aspects of human semantic processing. From the perspective of technical applications, there is no doubt that VLMs can perform tasks, such as visual question answering and image captioning, that are simply impossible for language-only models. This opens up the intriguing question of whether VLMs also learn text representations that model human language processing more accurately.
Regarding human behavioural responses, Pezzelle, Takmaz, and Fernández (2021) found promising evidence that VLM embeddings correlate with human similarity judgments more strongly than embeddings from language-only models when the judgments concern highly concrete concepts. As for brain responses, some recent works have explored the brain-modelling potential of VLMs using a variety of stimuli—including pictures Conwell et al. (2023a); Doerig et al. (2022); Wang et al. (2023), audio stories Tang et al. (2023), movies Tang et al. (2023); Dong and Toneva (2023); Subramaniam et al. (2024), and combinations of pictures and textual stimuli Oota et al. (2022); Nikolaus et al. (2024)—and a variety of experimental procedures, such as brain encoding Dong and Toneva (2023); Subramaniam et al. (2024); Oota et al. (2022), cross-encoding Tang et al. (2023) and cross-decoding Nikolaus et al. (2024). However, most of these studies focus on brain responses elicited in complex multimodal settings (e.g., watching movies or listening to long stories), as opposed to linguistic processing of single concepts.
Closer to our focus, Oota et al. (2022) used VLMs as the basis for encoding models of functional magnetic resonance imaging (fMRI) responses to concepts presented through written words and images simultaneously. We extend this work by additionally analysing fMRI responses in a purely linguistic condition and evaluating additional VLM architectures, while using a simpler method for measuring representational alignment. More concretely, we extract representations of concept words with a set of pre-trained VLMs, and apply Representational Similarity Analysis (RSA, Kriegeskorte, Mur, and Bandettini, 2008) to compute their alignment with human fMRI responses recorded while participants are viewing written concept words, either accompanied by pictures (picture condition) or embedded in sentences (sentence condition); see Figure 1 for illustration.
Since vision-language models largely differ in architecture and specific strategies employed to achieve multimodal integration, we identify three main families—contrastive VLMs, vision-language encoders, and generative VLMs—and include representatives from each in our experiments. In addition, we evaluate a set of language-only architectures that roughly match the language model component in the VLMs we consider.
In summary, we investigate two key research questions:
  1. (1)
    Are representations by current pretrained VLMs better models of brain activity involved in concept word processing than text-only language models? As mentioned above, human concept representations seem to reflect knowledge from different modalities. Since VLMs are trained to align input from visual and linguistic streams, we hypothesise they will exhibit an advantage over text-only language models in modelling brain activity during concept processing.
  2. (2)
    How does the brain alignment of different VLM families vary across conditions and brain networks? The experimental conditions we consider differ fundamentally in that the target concept words are presented within either a sentential or a visual context. The presence vs. absence of a sentential context is known to affect the way human brains respond to language Xu et al. (2005); similarly, the ability to process longer contexts is known to result in stronger capabilities at downstream tasks in DNN models. We conduct an extensive comparison of VLM architectures and families, considering three brain networks for each condition; this allows us to understand which specific aspects of brain language processing are better captured by each model.
Altogether, the present work makes the following contributions:
  • It provides, for the first time, a systematic and exhaustive investigation of the brain alignment achieved by transformer-based pretrained VLMs belonging to different model families;
  • It shows evidence that the highest brain alignment is consistently achieved by one of the VLMs (and not a language-only model), although not the same architecture across all conditions and brain networks;
  • It reveals that vision-language encoders, such as LXMERT Tan and Bansal (2019) and VisualBERT Li et al. (2019), tend to exhibit higher brain alignment than the more recent generative VLMs;
  • It demonstrates that the superior brain alignment achieved by vision-language encoders stems from learning novel semantic information and cannot simply be attributed to receiving additional visual input or to semantic information already present in their language-only backbone.
We open-source code and materials to reproduce our experiments at the following public GitHub repository: https://github.com/dmg-illc/concept-processing-brain-vlms.
2 Background
2.1 Use of deep neural networks to model human brain activity
DNN models trained on large amounts of data with task-optimized (or, alternatively, self-supervised) representation-learning algorithms currently underlie virtually all state-of-the-art AI systems. Such models have enabled unprecedented advances in the past decade across the fields of computer vision and natural language processing, as well as applications integrating the two Mitchell et al. (2019). Meanwhile, computational cognitive (neuro)science has made impressive progress using DNNs as model systems of human behaviour and its neural substrates. Traditional experimental approaches often rely on very specific theory-driven hypotheses, with accordingly designed controlled stimuli and contrasting conditions. On the other hand, DNNs provide a complementary approach to studying human information processing, by generating rich representational spaces to encode more complex naturalistic stimuli. In a relatively new line of studies, DNNs are employed as stimulus-computable models, whose internal representations and output behaviour are directly compared to human neural and behavioural responses measured when fed with similar input (Kriegeskorte, Mur, and Bandettini, 2008; Ma and Peters, 2020; Doerig et al., 2023).
Internal representations from DNNs have proven remarkably predictive of human brain activity in both visual and linguistic processing. In the domain of vision, early studies using convolutional neural networks (CNNs) optimised for object recognition found that representations from such models successfully predict fMRI (Güçlü and Van Gerven, 2015; Eickenberg et al., 2017), magnetoencephalography (MEG; Seeliger et al., 2018) and neural spiking (Yamins et al., 2014) responses to natural images. More recent work comparing CNNs to the larger vision-transformer architectures has shown that performance at predicting visual cortex activity is in fact comparable between models with more human-like architectural biases (e.g., convolutional processing mechanisms, as in CNNs) and models with more human-like training objectives (e.g., no direct category supervision, as in vision transformers), when training data is controlled for (Conwell et al., 2023b).
Regarding models of language, early work pioneered the use of corpus-derived and neural network-based static word embeddings as predictive models of brain activity in linguistic tasks such as word reading/processing (Mitchell et al., 2008), sentence reading (Pereira et al., 2018) and narrative understanding (Wehbe et al., 2014). However, a huge improvement in this domain has been achieved by employing internal states from LSTM- and, later, transformer-based language models (e.g., Jain and Huth 2018; Toneva and Wehbe 2019; see Karamolegkou, Abdou, and Søgaard 2023 for a review), whose ability to learn contextualised representations is more akin to the human online processing of linguistic stimuli in context. Most prominently, Schrimpf et al. (2021) compared a comprehensive range of neural network-based embeddings—including static GloVe embeddings (Pennington, Socher, and Manning, 2014), LSTMs (Hochreiter and Schmidhuber, 1997), BERT-based masked language models (Devlin et al., 2019) and GPT-like autoregressive language models (Radford et al., )—in their predictivity of human neural (functional magnetic resonance imaging and electrocorticography) and behavioural (reading time) data and found that generative, autoregressive architectures were most predictive of human cognitive signals in language processing.
The wide applicability of pre-trained DNN hidden state representations and derived model output measures highlights one of the major strengths of this class of models compared to other cognitive modelling paradigms: Models trained to capture a rich variety of phenomena are more likely to be useful for multiple down-stream tasks and also to overlap in interesting ways with similar general-purpose representations involved in human language processing (Antonello and Huth, 2024).
2.2 Early multimodal models of meaning representations
Despite successes in applying DNNs to neural and cognitive modelling across the visual and linguistic domains, most studies described above concern unimodal DNNs, encoding either visual or textual input but not both. However, earlier approaches in computational semantics have repeatedly found that models of lexical meaning that combine visual and textual information are better aligned with human word representations than their unimodal counterparts. Many studies in this line of work combined modalities by concatenating textual and perceptual vectors, often drawing on separate sources of data for building concept representations in either domain (Bruni et al., 2012; Kiela and Bottou, 2014; Kiela, Verő, and Clark, 2016; Derby et al., 2018; Davis et al., 2019); in parallel, there have been efforts to construct unified multimodal embeddings aimed at representing the latent information shared across textual and perceptual modalities (Lazaridou, Pham, and Baroni, 2015; Zablocki et al., 2018; Silberer and Lapata, 2012, 2014).
Overall, such multimodal distributional semantic models were shown to outperform text-only approaches at tasks requiring visual knowledge, as well as more general semantic benchmarks measuring, for example, correlations to human word similarity or relatedness judgments (e.g., WordSim-353, Finkelstein et al. 2001; SimLex999, Hill, Reichart, and Korhonen 2015; MEN, Bruni et al. 2012). Moreover, concatenated visio-linguistic embeddings also showed an advantage over text- and vision-only baselines in decoding brain activity related to concrete noun representations (Davis et al., 2019). More recent studies have evaluated representations extracted from fully end-to-end models trained to map between images and corresponding text captions and found that such models show improved performance over text models on sentence representation tasks (Kiela et al., 2018) as well as improved alignment to human word similarity and relatedness judgments (Merkx, Frank, and Ernestus, 2022). These findings beg the question of whether such emergent cross-/multimodal representations in end-to-end models are also well-aligned with brain activity involved in human concept processing.
2.3 Multimodal semantic representations in the era of vision-language transformers
In recent years, a powerful class of transformer-based VLMs have achieved outstanding performance on a variety of downstream AI tasks. After being pre-trained on matched visual and textual data (e.g., image-caption pairs), these models turn out particularly capable at tasks requiring integrating information from textual and visual inputs, such as visual question answering, image captioning, and visual dialogue (see Suglia, Konstas, and Lemon 2024 for a review). In addition to such broad applications, VLMs have also been evaluated on specific aspects of human-like semantic and situational understanding, such as compositionality (Thrush et al., 2022a), goal-based action selection (Pezzelle et al., 2020) and embodied simulation (Jones and Trott, 2024), as well as their capabilities to deal with particular linguistic phenomena, such as gradable adjectives (Pezzelle and Fernández, 2019) and other constructions (Parcalabescu et al., 2022a; Chen, Fernández, and Pezzelle, 2023). These studies indicate that, despite their success in downstream applications, VLMs often lag behind humans on such targeted linguistic evaluations, opening the question of how closely their semantic representations are truly aligned with human ones.
Beyond their applications in AI, VLMs have also been evaluated in their ability to capture information relevant for modelling human brain responses. A few works on visual processing have demonstrated the advantages of VLMs over vision-only models in matching human brain responses. For instance, Wang et al. (2023) showed that image embeddings from CLIP (a model trained to match embeddings of images and corresponding text captions in a joint space; Radford et al., 2021) are significantly more predictive of high-level visual cortex activity in picture viewing than models trained on simpler image-label pairs or text only.
Another line of work has used VLMs to model brain responses to more complex multimodal stimuli, such as movies, stories, and combinations of images and text, offering mixed evidence regarding multimodal models’ ability to learn information beyond that learnt by unimodal models. For example, Nikolaus et al. (2024) focussed on cross-decoding fMRI responses to pictures and their textual descriptions and reported that multimodal representations do not increase decoding performance compared to unimodal (vision-only and language-only) representations. Similarly, Dong and Toneva (2023) used a multimodal transformer (processing video frames, text, and audio) to model fMRI recordings from participants watching a popular TV show and found that vision-audio representations were more predictive of brain activity in language regions than audio-only and vision-only representations; however, the brain predictive power of the multimodal representations did not go beyond the combination of that of unimodal representations. This partially contrasts with the findings by Tang et al. (2023), who used representations from VLMs to train encoding models generalising between fMRI responses to movies and audio stories about the same contents, and showed that this encoding performance was significantly higher than that by unimodal representations (vision-only and language-only). Finally, Subramaniam et al. (2024) used representations from VLMs to predict stereo-electroencephalography (sEEG) responses to movies, and identified sites of multimodal brain integration by checking in which areas the VLM representations were more predictive than language-only, vision-only, and linearly-combined vision and language representations.
Moving to single-word processing, Pezzelle, Takmaz, and Fernández (2021) evaluated word representations extracted from several VLM architectures against human semantic judgments, including the word similarity and relatedness benchmarks mentioned in Section 2.2. VLM word representations aligned better with human judgments than representations from text-only models, although only on concrete word pairs. Furthermore, Zhuang, Fedorenko, and Andreas (2024) evaluated a set of language-only, vision-only, and vision-language architectures on several measures of word learning over model training, and found that visual information is beneficial only when small amounts of training data are available. However, none of these works involved brain responses.
Taken together, these works suggest that VLMs are capable—although not always more capable than unimodal models—to predict brain responses to multimodal stimuli and capture some cognitively relevant aspects of single-word representations. A natural next question is whether VLMs are better than language-only models at predicting brain responses from isolated concept words. This issue was partially addressed by Oota et al. (2022), who predicted fMRI responses to single concepts—presented through both images and text—with representations by unimodal (language-only and vision-only) models and VLMs against fMRI responses and found VisualBERT Li et al. (2019) to be more brain-predictive than the other architectures. We extend this work by including fMRI responses to purely textual stimuli and evaluating representations from additional, more recent VLMs.
3 Methods
To study the alignment between vision-language models (VLMs) and brain responses, we focus on a dataset of neural responses to concept words, which participants read either accompanied by pictures or within sentences (Experiment 1 in Pereira et al., 2018). We derive embeddings from the same set of concept words using a set of VLMs and language-only models and quantify their alignment with human responses in different brain networks with representational similarity analysis (RSA, Kriegeskorte, Mur, and Bandettini, 2008). Below we describe the data, the models used, and our experimental procedures in detail. An overview of our experimental setup is provided in Figure 1.
Figure 1: Overview of the experimental setup in the sentence (top) and picture (bottom) condition, including the stimuli used for embedding extraction. Each concept word embedding is averaged over the model’s hidden states across 6 instances of processing a concept word (contextualized by different sentences or pictures), as are the recorded fMRI responses. Best viewed in colour.
3.1 Data
We evaluated models’ alignment to brain responses on the dataset introduced by Pereira et al. (2018). Specifically, we focus on part of the data used in their Experiment 1, which consists of voxel-wise fMRI activations collected while participants are exposed to words representing specific concepts in different conditions. The dataset consists of 180 words, including different parts of speech (128 nouns, 22 verbs, 29 adjectives and adverbs, and 1 function word, see Appendix A for the full list).
We consider two experimental conditions: a language-only sentence condition and a multimodal picture condition. In the sentence condition, each word appears boldfaced in the context of a sentence that makes the relevant concept (i.e., the target sense of the word) salient. For each concept word, participants see six sentences, one at a time. In the picture condition, each word is presented together with an image illustrating the relevant concept. Again, for each concept word, participants are shown six different images, one at a time. 16 participants were scanned while viewing the words in these two conditions. In both cases, they are invited to think about the meaning of the target concept in the context where it appears, i.e., the accompanying sentence or image.
Pereira et al. (2018) preprocessed the fMRI data by estimating the response to each stimulus using a general linear model in which each stimulus presentation (sentence or word and picture) was modeled with a boxcar function convolved with the canonical hemodynamic response; throughout this paper, we refer to the resulting voxelwise beta estimates as voxel activations or brain responses. Voxel activations for each participant are averaged across the six presentations of the same word per condition. Hence, we have one averaged brain response measurement per concept, condition, and participant.
We further analyse subsets of this measurement from two functionally localised networks targeted by Pereira et al. (2018), involved in either linguistic or visual processing: the Language network (Fedorenko, Behr, and Kanwisher, 2011) and the Visual network (Power et al., 2011; Buckner, Andrews-Hanna, and Schacter, 2008). Anatomical regions of interest included as part of each of these functional networks are listed in Table 1. We present results considering bilateral activity for the visual network, while, for the language network, results are reported separately for the left hemisphere (LH Language) and the right hemisphere (RH Language), as done by Pereira et al. (2018).
Network Regions of Interest
Language - Posterior temporal lobe - Inferior frontal gyrus
- Anterior temporal lobe - Middle frontal gyrus
- Angular gyrus - Orbitan inferior frontal gyrus
Visual - Parahippocampal place area - Superior temporal sulcus
- Retrosplenial cortex - Fusiform face area
- Transverse occipital sulcus - Occipital face area
- Lateral occipital area - Extrastriate body area
Table 1: Brain regions of interest (ROIs) corresponding to our investigated functionally localised language and visual networks.
3.2 Models
For our experiments, we employ two main types of models: a set of VLMs trained on related visual and textual input and a set of language-only models trained exclusively on text. To make our evaluation as comprehensive as possible, we include multiple VLMs implementing vision-language integration with different strategies. We elaborate more on this point below (Section 3.2.1). Furthermore, we are interested in testing powerful models achieving high accuracy on downstream tasks to check whether their performance goes hand in hand with the human-likeness of their language representations. Therefore, we focus on VLMs that are widely known within the NLP community as they are—or were, when they were released—state-of-the-art in downstream applications, and use them off-the-shelf as pre-trained by their creators. Regarding the language-only models, we include architectures that either provide informative baselines (e.g., GloVe) or useful comparisons with specific VLMs (e.g., Llama3 is the language module of the VLM LLaVA). We outline the models’ main features below.
3.2.1 Vision-language models
VLMs can handle mixed inputs that include both text and images thanks to their pre-training on data from both modalities. In this sense, the pre-training data VLMs learn from is richer than the pre-training data language-only models are exposed to. A question we aim to investigate here is whether this richer input translates into representations that are more similar to those developed by humans, who are typically exposed to stimuli from different modalities when learning concepts.
While most VLMs are transformer-based, the modules making up their architecture and the way in which they achieve visuo-linguistic integration may vary. Based on these elements, it is possible to identify three main types or ‘families’ of VLMs, i.e., contrastive vision-language models, vision-language encoders, and generative vision-language models. A schematic overview of the three model families is provided in Figure 2.
Figure 2: Overview of the core components constituting the three families of VLMs described in Section 3.2.1. The diagrams exemplify the input provided to the models in the picture condition. The model outputs at the bottom of each diagram are the representations we compare against brain responses through representational similarity analysis. Best viewed in colour.
Contrastive VLMs
The contrastive vision-language models we experiment with are CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These models encode image and text separately with two dedicated transformer-based modules and are pre-trained with a contrastive loss, which maximises the similarity between the image and text embeddings of matching image-text pairs. The CLIP implementation we deploy has the transformer-based ViT (Dosovitskiy et al., 2021) as image encoder and a language transformer as text encoder. As for ALIGN, it uses the convolutional neural network EfficientNet (Tan and Le, 2019) to encode images and the transformer-based BERT (Devlin et al., 2019) to encode text.
Vision-language encoders
The vision-language encoders employed in our experiments are VisualBERT (Li et al., 2019) and LXMERT (Tan and Bansal, 2019). Their distinctive feature is a specific module that uses attention mechanisms to learn relations between visual features extracted with an object detector—in the implementation we use, the convolutional-neural-network-based Faster RCNN (Ren et al., 2015)—and text embeddings. In VisualBERT, language processing and multimodal integration are performed by the same BERT-based module. In contrast, LXMERT has two transformer-based unimodal encoders (with the language encoder being BERT-based) and an additional transformer-based module to perform multimodal integration. Both models build text representations that incorporate information from the visual input and are, in this sense, multimodal.
Generative VLMs
For this family of VLMs, we consider IDEFICS2 (Laurençon et al., 2024) and LLaVA NeXT (Liu et al., 2024). Their core module is a transformer-based decoder-only large language model—Mistral (Jiang et al., 2023) in IDEFICS2 and Llama3 (Grattafiori et al., 2024) in LLaVA NeXT—that is pre-trained autoregressively and can therefore generate textual outputs based on the provided textual prompts and visual inputs. In order for their outputs to be informed by the visual inputs, these models are equipped with transformer-based visual encoders—SigLIP-SO400M (Zhai et al., 2023) in IDEFICS2 and ViT (Dosovitskiy et al., 2021) in LLaVA NeXT—and projection modules which learn to attend to the visual information useful for text generation.
3.2.2 Language-only models
We experiment with one encoder-only language model—BERT—and two generative large language models—Mistral, Llama3. In addition, we include a simpler distributional semantic model—GloVe—for reference.
BERT (Devlin et al., 2019) is a transformer-based encoder-only model pretrained with masked language modelling and next-sentence prediction. This language-only model forms the basis of the language components of the contrastive VLM ALIGN and the vision-language encoders VisualBERT and LXMERT. Note that, while VisualBERT is initialised with BERT weights, LXMERT and ALIGN train their BERT-based modules from scratch.
Mistral (Jiang et al., 2023) and Llama3 (Grattafiori et al., 2024) are decoder-only large language models (LLMs) pre-trained autoregressively. Their key feature is that they can generate natural-language outputs based on the input prompts. We choose these specific LLMs as Mistral is the language model used in IDEFICS2 and Llama3 is used in LLaVA NeXT.
Finally, GloVe (Pennington, Socher, and Manning, 2014) is a type-based distributional model that learns word embeddings based on co-occurrence patterns. In contrast to all the other models we deploy, GloVe representations are not contextualised. Nevertheless, they have been shown to moderately correlate with brain responses as described in Section 1. We therefore include them, as they still provide a useful reference.
3.3 Procedure
To extract model representations that we can compare to the fMRI responses, we feed the models with the same stimuli presented to the participants in the experiments by Pereira et al. (2018). Different models require slightly different inputs, as we explain in detail in this section.
Sentence condition
The stimuli passed to the tested models in this condition are the sentences read by participants during the sentence-condition part of Experiment 1 in Pereira et al. (2018). After extracting model representations for the entire sentences, we isolate the parts of the representations corresponding to the tokens of the target concept word.
Regarding the language-only models, for the transformer-based models (BERT, Mistral, and Llama3), we provide each stimulus sentence as input to the model and extract the representations of the target word from the models’ hidden states. [^1] For all models, if the target word is made up of several tokens, we average over token representations. As for GLoVe, given that it provides static (type-based) decontextualised representations, we simply select the pre-trained GloVe embedding corresponding to each target word.
Regarding the contrastive VLMs (CLIP and ALIGN), we pass sentences to their language encoder, and extract the target word representations as in the transformer-based language models. The procedure for VisualBERT is similar. Due to technical reasons, LXMERT requires including some visual input in addition to text. To make this input as neutral as possible, together with each sentence, we pass to the model an image made up of random noise. [^2] We create a different random noise image for each sentence. These images are included in the supplementary materials linked to our public GitHub repository. As for generative models, sentences are passed to the LLM-based language decoder. The target word representations are extracted from the hidden states of the language encoder in the case of LXMERT and VisualBERT, while for the generative VLMs (LLaVA NeXT and IDEFICS2), word representations are extracted from the language modelling decoder.
To obtain a final representation for each concept word that can be compared to brain activations, we average the representations extracted from the 6 sentences where the word appears (except for GloVe, where sentences are not considered). Furthermore, for the models where the architecture allows it (all except GloVe), we repeat the procedure described above for each model layer.
Picture condition
In the Picture condition, participants see each word together with an image. For the VLMs, both the target word and each image accompanying it are fed to the models. When using contrastive models, we integrate the image and text embeddings by taking their element-wise sum. For the VL encoders, we extract hidden states from all text-specific layers and from the cross-modal layers. In generative VLMs, we consider the hidden states from all layers of the LLM module. To obtain a final representation for each concept word that can be compared to brain activations, we then average the extracted representations across the six images per word.
Naturally, for the language-only models, only the target word is provided as input. The obtained representations are therefore not linguistically contextualised in this case, in contrast to the sentence condition.
3.4 Evaluation
We evaluate the models’ alignment to fMRI data with Representational Similarity Analysis (RSA, Kriegeskorte, Mur, and Bandettini, 2008). This method quantifies the similarity between representational geometries by comparing the pairwise distances between stimulus representations in two representational spaces, across a joint set of stimuli. More specifically, in our experiments we compare cosine distances between stimulus embeddings derived from the models and cosine distances between fMRI responses to different stimuli. We populate two representational dissimilarity matrices (RDMs) with brain- and model-derived pairwise distances and then quantify their representational similarity (RS) using Spearman’s correlation coefficient over the upper triangle of these matrices (i.e., the unique set of pairwise distances in each space). A significant positive correlation between RDMs indicates that there is alignment between brain-derived and model-derived representational geometries.
We compute representational similarity against brain RDMs averaged across single participants. For each model with multiple layers, RSA is computed separately for model representations from each layer. When drawing comparisons between models, we consider the best (i.e., most aligned) layer for each model. We verify that differences in brain alignment between models are statistically significant by applying a Fisher transformation to all the unique pairs of Spearman correlations and calculating the $p$-value associated with the difference between the two $z$-scores. To control for false positives due to multiple comparisons (45 per brain network in each condition), $p$-values are Bonferroni-corrected with $\alpha$ = 0.05. We report the complete set of $p$-values in Appendix B.
To obtain a random baseline for the representational similarity, we computed alignment against shuffled brain responses, i.e., a condition where each concept was associated with a randomly chosen fMRI response from a non-matching concept.
4 Results
Representational similarity analysis (RSA) reveals many interesting patterns in model representations’ alignment to fMRI responses. These results are visualised in Figures 3 (sentence condition) and 4 (picture condition), which report representational similarity from the most brain-aligned model layer in the cases where we extract representations from multiple model layers. Layer-wise RSA values for all models are provided in Appendix C. In general, we observe that the best layer for each model is rarely the same across brain networks and experimental conditions. Nevertheless, there are a few interesting correspondences that we highlight in the following subsections.
Figure 3: RSA results for the sentence condition. Spearman $\rho$ indicates the correlation between model-derived concept representations and fMRI responses in the left-hemisphere (LH) language network, right-hemisphere (RH) language network, and visual network. The numbers in brackets after model names indicate the layer from which representations were extracted. Best viewed in colour.
Figure 4: RSA results for the picture condition. Spearman $\rho$ indicates the correlation between model-derived concept representations and fMRI responses in the left-hemisphere (LH) language network, right-hemisphere (RH) language network, and visual network. The numbers in brackets after model names indicate the layer from which representations were extracted. Best viewed in colour.
4.1 Sentence condition
As displayed in Figure 3, in the sentence condition, all models exhibit low to moderate positive correlations with brain responses from all three networks, except for ALIGN in the RH (right-hemisphere) language network. All these correlations are statistically significantly different from 0, which coincides with the random baseline computed by shuffling brain responses. A closer examination of models’ brain alignment in each of the brain networks reveals the following insights.
In the LH language network, the VL encoder VisualBERT is statistically significantly more brain-aligned than all the other models ($\rho$ = 0.10) except for LXMERT ($\rho$ = 0.08). Another interesting result is that GloVe’s correlation with fMRI responses is not statistically significantly different from that of the more advanced models—both multimodal and unimodal—CLIP, LXMERT, IDEFICS2 and Mistral. Comparisons between VLMs and their language-only counterparts reveal that both VisualBERT and LXMERT are significantly more brain-aligned than BERT, while LLaVA and IDEFICS2 are not different from their language decoders Llama3 and Mistral.
In the RH language network, the most brain-aligned models are IDEFICS2 and Mistral ($\rho$ = 0.08 for both), which, however, are not statistically significantly superior to CLIP, VisualBERT, LXMERT and LLaVA. The lowest alignment is observed for the word-embedding model GloVe ($\rho$ = 0.01) and the contrastive model ALIGN, which exhibits a negative correlation ($\rho$ = -0.06). This pattern—high alignment for models that can handle long contexts and low alignment for models that privilege single-word representations—fits well with the notion that the right brain hemisphere is particularly involved in modelling contextual aspects of language Xu et al. (2005).
In the visual network, we observe the highest correlations for VisualBERT and IDEFICS2 ($\rho$ = 0.09 for both), which are statistically significantly more brain-aligned than all other models except for LXMERT and Mistral. Regarding the remaining comparisons between VLMs and language-only counterparts, LLaVA’s advantage over Llama3 is not statistically significant, while both VisualBERT and LXMERT are statistically significantly more brain-aligned than BERT.
Overall, our results in this condition indicate that the VLMs IDEFICS2 and VisualBERT tend to exhibit the strongest brain alignment across all brain areas. However, IDEFICS2’s alignment to brain responses in both language areas is not statistically significantly different from Mistral’s; this, together with the fact that both models exhibit the highest brain alignment in the same or neighbouring layers, suggests that their brain-modelling properties in these networks are highly similar. Considering VLM families, VL encoders are statistically significantly superior to contrastive VLMs in the LH language network and in the visual one—but not in the RH language network—and to generative VLMs in the LH language network. Lastly, VL encoders consistently outperform their language-only counterpart BERT.
4.2 Picture condition
RSA results for the picture condition are displayed in Figure 4. In general, we observe higher correlations than in the sentence condition, which can be attributed to higher signal (i.e., higher inter-participant similarities) in the fMRI responses. As in the sentence condition, all the correlations are statistically significantly different from the random baseline 0.
In the LH language network, both contrastive VLMs and VL encoders exhibit moderate correlations with brain responses ($\rho$ > 0.3), which are all statistically significantly stronger than those achieved by generative VLMs and language-only models. In the RH language network, VisualBERT outperforms all other models by a significant margin ($\rho$ = 0.37) and contrastive VLMs and VL encoders are significantly more brain-aligned than the other model families. Correlations in the visual network exhibit a pattern similar to the RH language network, with VisualBERT outperforming all other models ($\rho$ = 0.48) and CLIP, ALIGN and LXMERT exhibiting similar brain correlations (0.25 < $\rho$ < 0.29). The similarities between these two networks (RH language and visual) are also evident within models, in the sense that, for LXMERT and VisualBERT, the most brain-aligned layers are the same across the visual network and the RH language one.
Two interesting findings common across all three brain networks are that all VLMs are statistically significantly more brain-aligned than their unimodal counterparts, and that in multiple models (LXMERT, IDEFICS2, LLaVA, Mistral, and Llama3) the most brain-aligned layers are among the latest.
Lastly, another trend common to all brain networks is that GloVe exhibits higher correlations than all other more powerful language-only models.  [^3] The only exception is the visual network, where the difference between GloVe and BERT is not statistically significant. As we further elaborate in Section 6, a potential reason for this may be that Pereira et al. (2018) selected the concept stimuli to be included in their experiment with a procedure based specifically on GloVe representations.
5 Ablation Studies
To complement the findings provided by RSA, we conduct two ablation analyses aimed at better understanding what drives VLMs’ brain alignment in both experimental conditions. More specifically, we investigate the role played by the semantic information provided by language-only modules in the sentence condition and by the input images in the picture condition. We detail the goals and procedures behind both analyses in the following.
5.1 Semantic information shared with language modules in the sentence condition
The first question that we explore to delve deeper into our initial results is: How much of the VLMs’ brain alignment in the sentence condition can be attributed to semantic information already present in their language encoder prior to any multimodal training? To investigate this, we conduct a partial correlation analysis aimed at removing from VLMs’ representational spaces the information shared with LLMs’ representational spaces. We focus on a subset of models whose architecture is highly similar to that of the language-only counterpart, i.e., VisualBERT, LXMERT, LLaVA and IDEFICS2. Note that, while VisualBERT, LLaVA and IDEFICS2 were initialised with pretrained weights from their language-only counterparts, ALIGN and LXMERT were trained multimodally from scratch.
More formally, partial correlations are computed as follows. Consider the RDM from a VLM $y$, the RDM from its language-only counterpart $x$, and the residuals $r_{i}=y_{i}-\hat{y}_{i}$ from the linear regression with equation $\hat{y}_{i}=a+bx_{i}$. The partial correlation achieved by the VLM is defined as the Spearman correlation $\rho(r_{i},z_{i})$, where $z$ represents the RDM of the brain responses, and provides an indication of the VLM’s brain alignment once the representational information shared with the language-only counterpart is ablated.
We display initial VLM correlations with brain responses along with partial correlations in Figure 5. If the brain alignment of a VLM is driven by its language-only component, the partial correlation should be significantly lower than the initial correlation. In contrast, the absence of such a difference indicates that the semantic knowledge learnt by the VLM is independent from that provided by its language-only module.
In the LH language network, all differences between partial and initial correlations—except for IDEFICS2 in language LH—are not statistically significant, suggesting that the brain alignment achieved by these models cannot be attributed to semantic information already present in the language-only modules, but to different knowledge acquired during multimodal pretraining. In the RH language network, differences between initial and partial correlations are statistically significant for VisualBERT, LLaVA and IDEFICS2, which, crucially, are the models initialised with pretrained weights from their language-only modules. Another interesting observation is that, while IDEFICS2’s brain correlation becomes null after the ablation, VisualBERT and LLaVA still retain part of their alignment ($\rho$ = 0.04 and $\rho$ = 0.03, respectively). Lastly, results from the visual network reveal diverging patterns between generative VLMs and the other VLMs: The differences between initial and partial correlations are statistically significant for the former, while they are not for the latter. Curiously, this suggests that part of the information relevant for alignment with visual brain responses was already present in Mistral and Llama3 before any vision-language training.
Figure 5: Spearman correlations quantifying initial VLMs’ alignment with fMRI responses in the sentence condition vs. the alignment obtained after ablating the information shared with the language-only counterparts. Asterisks indicate whether differences between initial correlations and partial ones are statistically significant ($p$-values Bonferroni-corrected with $\alpha$ = 0.05 and four comparisons). Best viewed in colour.
5.2 Visual information in the picture condition
Figure 6: Results from the ablation study where we pass only concept words to both VLMs and language-only models. For each brain network, we show the Spearman correlations resulting from RSA, indicating the alignment between models and fMRI responses from the picture condition. Numbers in brackets indicate the layers from which representations are extracted. Best viewed in colour.
In the picture condition, we documented a systematic advantage of VLMs over their language-only counterparts (see Section 4.2). However, a potential confound could be that VLMs have access to additional contextual information—the picture—that is available to the human participants but not to language-only models. In this sense, their superior alignment with human responses could stem from an uneven comparison rather than indicate that they capture additional semantic information.
To shed more light on this issue, we conduct an ablation study where we pass only the concept word (without pictures [^4] As mentioned above, this is not possible for LXMERT, as the model requires a visual input. Therefore, similarly to what we do in the sentence condition, we pass a random-noise image.) to the VLMs, so that the input they receive matches that provided to language-only models. We then recompute RSA against human fMRI responses from the picture condition, following the same procedures employed in the main experiment. The resulting Spearman correlations are provided in Figure 6. We assess the statistical significance of differences between each pair of correlations per network as in the main experiment.
This ablation study reveals several interesting trends. First, comparing VLM’s alignment when the pictures are provided vs. when they are not shows that correlations decrease across all brain areas, sometimes dramatically, for CLIP, ALIGN, VisualBERT, and LLaVA (see Figure 7 in Appendix D to facilitate comparisons). On the other hand, the variations observed for LXMERT and IDEFICS2 are minimal, suggesting that their initial brain alignment is not merely a byproduct of the additional context provided by the images.
Despite changes in model rankings, the most brain-aligned architectures in all brain networks remain multimodal. More specifically, LXMERT is statistically significantly more brain-aligned than all other models across all three networks, and VisualBERT statistically significantly outperforms all language-only models in the LH language network and in the visual one.
Considering VLMs and their language-only counterparts, LXMERT and VisualBERT consistently outperform BERT with statistically significant improvements, while ALIGN is less brain-aligned than BERT across all brain networks. As for the generative VLMs, IDEFICS2 remains statistically significantly more brain aligned than Mistral and LLaVA is significantly less brain-aligned than LLama3 in the RH language network and in the visual one.
Overall, this analysis sheds light on the extent to which VLMs’ brain alignment is driven by the input images, suggesting that, while some architectures rely heavily on them, others yield strong brain correlations even without meaningful visual input.
6 Discussion
A key notion to consider when interpreting our results is the fundamental difference between the sentence experimental condition and the picture one. Indeed, while both conditions focus on single-word semantics, the processes at play when understanding them within a textual vs. visual context may largely differ—for both humans and machines. From the perspective of human cognition, it has been repeatedly shown that processing isolated words as opposed to words presented within a sentential or narrative context results in brain responses that differ in terms of both temporal and spatial dynamics (e.g., Xu et al., 2005; Deniz et al., 2023). From a model viewpoint, there are substantial differences (in terms of architecture, model size, amount of training data, etc.) between models trained to output word embeddings vs. models trained to process longer text sequences and output contextualised word representations. In the following, we relate these considerations to the patterns and trends in models’ brain alignment revealed by our results.
6.1 Differences between language-only models
The differences between experimental conditions are interestingly reflected in two main patterns observed for language-only models. First, GloVe—the only word embedding model used in our study—proves to be consistently more brain-aligned than the other language-only models trained to handle longer text sequences in the picture condition, where words are presented in isolation (at least in terms of linguistic context), but not as systematically in the sentence condition, where there is additional linguistic context. Second, in the sentence condition, GloVe remains more brain-aligned than the other language-only models in the LH language network, but not in the RH one, where transformer-based models, and in particular the autoregressive one Mistral, achieve higher correlations with brain responses.
On the one hand, these findings align well with extant neuroscientific evidence showing that the left hemisphere is more specialised in the semantic aspects of language processing, while the right one becomes more involved as the contextual complexity of the text increases Xu et al. (2005). On the other hand, the same results may appear in contrast with previous computational studies documenting higher brain alignment for transformer-based models than word embedding ones (e.g., Schrimpf et al., 2021). However, previous and current findings can be reconciled in light of the differences in the experimental conditions. Namely, previous studies finding an advantage for transformer-based models focussed on brain responses to sentences or longer narratives, making them more comparable to our sentence condition—where an advantage for transformer-based models was indeed documented at least in the RH language network—but not to the picture condition, where such an advantage was not observed.
One remaining open question is why, in the sentence condition, we found that GloVe outperforms BERT, Mistral, and Llama3 in the LH language network. We speculate that this may be due to the procedure Pereira et al. (2018) employed to select the concept stimuli, which was based on spectral clustering of GloVe semantic vectors. Additional reasons for GloVe’s success may be linked to the pipeline used to collect and analyse the fMRI responses. More specifically, the entire experimental pipeline—from explicitly asking participants to focus on the meaning of boldfaced target concepts to analysing the fMRI responses only for those concepts—remains focused on isolated words, even in the sentence condition. Therefore, it is possible that transformer-based models have an advantage at the sentence level, but this was not detected given the focus of our analyses on single concept words.
6.2 Differences between VLM families
Considering the VLM families, an interesting finding was that VL encoders tend to be more brain-aligned than contrastive VLMs in the sentence condition, while contrastive VLMs are competitive with VL encoders in the picture condition. A possible explanation for this result is that contrastive VLMs are effective at capturing object-word correspondences—a scenario quite akin to the picture condition—but struggle to accurately represent more complex relations between entities Thrush et al. (2022b); Liu, Emerson, and Collier (2023); Parcalabescu et al. (2022b); Hendricks and Nematzadeh (2021). This challenge may, therefore, have resulted in lower performance compared to the VL encoders.
Regarding generative VLMs, the brain correlations they exhibit in the picture condition are lower than those by contrastive VLMs and VL encoders. At first glance, this result may seem counterintuitive, given the superior performance of generative VLMs in downstream tasks. However, the trend can be explained considering that autoregressive models have been shown to privilege production over representation as a result of their pretraining on next-word prediction BehnamGhader et al. (2024); Muennighoff et al. (2024); Springer et al. (2024). In this sense, using them in a scenario where a single word has to be encoded is a use case quite far from the tasks they were originally designed for.
Despite their limitations in the picture condition, it is worth highlighting that generative VLMs remain competitive with VL encoders in the sentence condition, especially in the RH language network. This can be explained by considerations analogous to those mentioned above for language-only models—namely, it stands to reason that generative models are better aligned with brain activity in a language network particularly involved in processing contextual aspects, given their ability to handle long contexts.
6.3 VLMs vs. their unimodal counterparts
Our results revealed an overall advantage of VLMs over language-only models, but the trend is not systematic across all VLM-language-only-model pairs and brain networks. In the picture condition, all VLMs were more brain-aligned than their unimodal counterparts, suggesting that their ability to process visual information is beneficial for modelling brain-relevant semantic aspects. In addition, the ablation study where we reassessed their brain alignment providing only the concept words suggests that an advantage attributable to multimodal pretraining is systematically present for LXMERT, IDEFICS and—in two of the three brain networks—VisualBERT
In the sentence condition, where all models received the same textual inputs, the picture emerging from the comparisons between VLMs and their counterparts is more nuanced, with VisualBERT and LXMERT outperforming BERT in the LH language network and in the visual one, and IDEFICS2, LLaVA and ALIGN never significantly outperforming Mistral, Llama3 and BERT, respectively. The analysis of partial correlations for the sentence condition further complements these findings, showing that the semantic contribution provided by ALIGN and VisualBERT is independent from that of their language encoder, while patterns were mixed for the generative VLMs, suggesting that, overall, they share more semantic information with their language decoders.
Taken together, our findings indicate that VL encoders exhibit a consistent advantage over BERT across experimental conditions in the majority of brain networks. This provides evidence that these models have learnt language representations that reflect brain-relevant semantic knowledge not captured by their unimodal counterparts. However, this trend is not systematic for all VLMs, suggesting that multimodal training is not per se enough for a model to learn more brain-like text representations.
7 Conclusion
The present study provides a broad investigation of the brain alignment achieved by vision-language models from different families, drawing meaningful comparisons with language-only models and considering two experimental conditions and three brain networks.
Our results from the sentence condition revealed that VLMs have an advantage over language-only models only on the LH language network, while still exhibiting competitive brain alignment in the remaining networks. This pattern was confirmed by the ablation study where we removed from VLMs’ representational space the knowledge shared with language-only models: While, in Language LH, VLMs display an advantage due to capturing multimodal information, alignment in the other areas is mostly driven by strictly linguistic information.
In the picture condition, we found that, in each brain network, there was at least one VLM outperforming all language-only models by a significant margin. This trend remained robust even in the ablation study where we provided the same input to both VLMs and language-only models, suggesting that the initial advantage we observed is not attributable solely to the information provided by the image, but to additional semantic information learnt during pretraining.
Overall, our findings indicate that VLMs’ advantage over language-only models in brain modelling is consistent in the left-hemisphere language network, while, in the others, it is substantial only when visual information is made salient to human participants. To conclude, our study provides a meaningful contribution to our understanding of the conceptual information learnt by multimodal and unimodal models by quantitatively evaluating its similarity with human brain responses. A question that remains open for future work is whether these findings generalise to situations where human participants are presented with more naturalistic stimuli.
[…]
Concepts from the Pereira Dataset
The full list of concept words from the Pereira dataset (Pereira et al., 2018) is the following:
Ability, Accomplished, Angry, Apartment, Applause, Argument, Argumentatively, Art, Attitude, Bag, Ball, Bar, Bear, Beat, Bed, Beer, Big, Bird, Blood, Body, Brain, Broken, Building, Burn, Business, Camera, Carefully, Challenge, Charity, Charming, Clothes, Cockroach, Code, Collection, Computer, Construction, Cook, Counting, Crazy, Damage, Dance, Dangerous, Deceive, Dedication, Deliberately, Delivery, Dessert, Device, Dig, Dinner, Disease, Dissolve, Disturb, Do, Doctor, Dog, Dressing, Driver, Economy, Election, Electron, Elegance, Emotion, Emotionally, Engine, Event, Experiment, Extremely, Feeling, Fight, Fish, Flow, Food, Garbage, Gold, Great, Gun, Hair, Help, Hurting, Ignorance, Illness, Impress, Invention, Investigation, Invisible, Job, Jungle, Kindness, King, Lady, Land, Laugh, Law, Left, Level, Liar, Light, Magic, Marriage, Material, Mathematical, Mechanism, Medication, Money, Mountain, Movement, Movie, Music, Nation, News, Noise, Obligation, Pain, Personality, Philosophy, Picture, Pig, Plan, Plant, Play, Pleasure, Poor, Prison, Professional, Protection, Quality, Reaction, Read, Relationship, Religious, Residence, Road, Sad, Science, Seafood, Sell, Sew, Sexy, Shape, Ship, Show, Sign, Silly, Sin, Skin, Smart, Smiling, Solution, Soul, Sound, Spoke, Star, Student, Stupid, Successful, Sugar, Suspect, Table, Taste, Team, Texture, Time, Tool, Toy, Tree, Trial, Tried, Typical, Unaware, Usable, Useless, Vacation, War, Wash, Weak, Wear, Weather, Willingly, Word.
[…]
Comparisons Between Models Table 2 provides all the ‘raw’ (i.e., not yet Bonferroni corrected) $p$-values we calculate in the main experiment to assess whether pairwise differences between models are statistically significant.
Model Pair LH Language RH Language Visual
Sent. Pic. Sent. Pic. Sent. Pic.
CLIP - ALIGN 0.001 0.014 0.0 0.976 0.051 0.0
CLIP - LXMERT 0.0 0.0 0.614 0.365 0.0 0.229
CLIP - VisualBERT 0.0 0.0 0.269 0.0 0.0 0.0
CLIP - IDEFICS2 0.745 0.0 0.011 0.0 0.0 0.0
CLIP - LLaVA 0.0 0.0 0.64 0.0 0.013 0.0
CLIP - Mistral 0.567 0.0 0.004 0.0 0.0 0.0
CLIP - Llama3 0.0 0.0 0.143 0.0 0.555 0.0
CLIP - BERT 0.001 0.0 0.171 0.0 0.234 0.0
CLIP - GloVe 0.097 0.0 0.0 0.0 0.001 0.0
ALIGN - LXMERT 0.0 0.004 0.0 0.382 0.0 0.0
ALIGN - VisualBERT 0.0 0.005 0.0 0.0 0.0 0.0
ALIGN - IDEFICS2 0.002 0.0 0.0 0.0 0.0 0.0
ALIGN - LLaVA 0.324 0.0 0.0 0.0 0.0 0.0
ALIGN - Mistral 0.004 0.0 0.0 0.0 0.0 0.0
ALIGN - Llama3 0.078 0.0 0.0 0.0 0.011 0.0
ALIGN - BERT 0.823 0.0 0.0 0.0 0.447 0.0
ALIGN - GloVe 0.0 0.0 0.0 0.0 0.147 0.0
LXMERT - VisualBERT 0.155 0.977 0.107 0.0 0.037 0.0
LXMERT - IDEFICS2 0.0 0.0 0.002 0.0 0.135 0.0
LXMERT - LLaVA 0.0 0.0 0.331 0.0 0.033 0.0
LXMERT - Mistral 0.0 0.0 0.001 0.0 0.982 0.0
LXMERT - Llama3 0.0 0.0 0.337 0.0 0.0 0.0
LXMERT - BERT 0.0 0.0 0.387 0.0 0.0 0.0
LXMERT - GloVe 0.008 0.0 0.0 0.0 0.0 0.0
VisualBERT - IDEFICS2 0.0 0.0 0.147 0.0 0.552 0.0
VisualBERT - LLaVA 0.0 0.0 0.524 0.0 0.0 0.0
VisualBERT - Mistral 0.0 0.0 0.082 0.0 0.035 0.0
VisualBERT - Llama3 0.0 0.0 0.01 0.0 0.0 0.0
VisualBERT - BERT 0.0 0.0 0.013 0.0 0.0 0.0
VisualBERT - GloVe 0.0 0.0 0.0 0.0 0.0 0.0
IDEFICS2 - LLaVA 0.0 0.01 0.037 0.0 0.0 0.0
IDEFICS2 - Mistral 0.805 0.0 0.774 0.0 0.129 0.0
IDEFICS2 - Llama3 0.0 0.0 0.0 0.0 0.0 0.0
IDEFICS2 - BERT 0.004 0.0 0.0 0.001 0.0 0.0
IDEFICS2 - GloVe 0.047 0.0 0.0 0.001 0.0 0.063
LLaVA - Mistral 0.0 0.0 0.018 0.0 0.035 0.0
LLaVA - Llama3 0.438 0.0 0.053 0.0 0.059 0.0
LLaVA - BERT 0.226 0.0 0.066 0.0 0.0 0.0
LLaVA - GloVe 0.0 0.0 0.0 0.0 0.0 0.001
Mistral - Llama3 0.0 0.0 0.0 0.916 0.0 0.001
Mistral - BERT 0.008 0.0 0.0 0.0 0.0 0.002
Mistral - GloVe 0.026 0.0 0.0 0.0 0.0 0.0
Llama3 - BERT 0.047 0.0 0.925 0.0 0.075 0.819
Llama3 - GloVe 0.0 0.0 0.0 0.0 0.0 0.0
BERT - GloVe 0.0 0.072 0.0 0.0 0.027 0.0
Table 2: $p$-values associated with all pairwise model comparisons in the main experiment before Bonferroni corrections.
[…]
Layer-wise RSA results
We provide layer-wise RSA results for the sentence condition in Tables 3, 6, 4, 7, 5 and 8. Layer-wise results for the picture condition are reported in Tables 9, 10 and 11. In each of these tables, we boldface the highest correlation value (which was identified before rounding to the second decimal) and check whether it is statistically significantly different from the correlation achieved by each of the other layers by performing pairwise comparisons as described in the main paper. $p$-values are Bonferroni-corrected, with the number of comparisons amounting to $1-\#layers$ for each model.
Layer CLIP ALIGN LXMERT VisualBERT BERT
1 -0.06 (0.0)* -0.0 (0.58)* 0.07 (0.0) 0.1 (0.0) 0.01 (0.13)
2 0.05 (0.0) -0.01 (0.22)* 0.07 (0.0) 0.08 (0.0) 0.02 (0.0)
3 0.03 (0.0) 0.01 (0.13) 0.07 (0.0) 0.07 (0.0) 0.02 (0.0)
4 0.02 (0.0)* 0.02 (0.01) 0.08 (0.0) 0.07 (0.0) 0.02 (0.01)
5 0.01 (0.43)* 0.01 (0.2) 0.06 (0.0) 0.07 (0.0) 0.02 (0.03)
6 -0.0 (0.64)* -0.0 (0.9) 0.05 (0.0)* 0.07 (0.0)* 0.02 (0.01)
7 0.01 (0.13)* 0.0 (0.81) 0.04 (0.0)* 0.06 (0.0)* 0.02 (0.05)
8 0.0 (0.68)* 0.0 (0.81) 0.03 (0.0)* 0.06 (0.0)* 0.01 (0.3)
9 0.01 (0.24)* 0.0 (0.63) 0.04 (0.0)* 0.05 (0.0)* -0.0 (0.58)*
10 -0.0 (0.66)* 0.0 (0.74) 0.04 (0.0)* 0.03 (0.0)* -0.03 (0.0)*
11 0.01 (0.12)* -0.0 (0.97) 0.03 (0.0)* 0.01 (0.22)* -0.05 (0.0)*
12 0.01 (0.1)* 0.0 (0.69) -0.03 (0.0)* -0.0 (0.57)* -0.05 (0.0)*
13 0.05 (0.0) 0.02 (0.02) -0.08 (0.0)* 0.0 (0.55)* -0.08 (0.0)*
14 - - 0.03 (0.0)* - -
Table 3: Layer-wise RSA values (Spearman correlations) for the sentence condition and LH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer CLIP ALIGN LXMERT VisualBERT BERT
1 -0.01 (0.38)* -0.08 (0.0)* 0.02 (0.0)* 0.06 (0.0) 0.02 (0.04)*
2 0.03 (0.0)* -0.08 (0.0)* 0.03 (0.0)* 0.04 (0.0) 0.02 (0.01)*
3 0.02 (0.0)* -0.06 (0.0) 0.04 (0.0) 0.06 (0.0) 0.03 (0.0)
4 0.02 (0.0)* -0.07 (0.0) 0.04 (0.0) 0.06 (0.0) 0.03 (0.0)
5 0.01 (0.13)* -0.08 (0.0) 0.04 (0.0) 0.06 (0.0) 0.04 (0.0)
6 -0.0 (0.89)* -0.07 (0.0) 0.03 (0.0) 0.06 (0.0) 0.03 (0.0)
7 0.03 (0.0)* -0.06 (0.0) 0.04 (0.0) 0.06 (0.0) 0.03 (0.0)
8 0.03 (0.0)* -0.06 (0.0) 0.04 (0.0) 0.06 (0.0) 0.03 (0.0)
9 0.04 (0.0) -0.07 (0.0) 0.02 (0.0)* 0.04 (0.0) 0.03 (0.0)
10 0.04 (0.0) -0.07 (0.0) 0.03 (0.0)* 0.04 (0.0)* 0.01 (0.09)*
11 0.06 (0.0) -0.07 (0.0) 0.03 (0.0) 0.03 (0.0)* 0.01 (0.28)*
12 0.03 (0.0) -0.06 (0.0) -0.01 (0.51)* 0.03 (0.0)* -0.0 (0.9)*
13 0.02 (0.01)* -0.06 (0.0) -0.04 (0.0)* 0.03 (0.0)* -0.01 (0.5)*
14 - - 0.05 (0.0) - -
Table 4: Layer-wise RSA values (Spearman correlations) for the sentence condition and RH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer CLIP ALIGN LXMERT VisualBERT BERT
1 -0.04 (0.0)* 0.0 (0.76) 0.06 (0.0) 0.09 (0.0) 0.02 (0.01)
2 0.04 (0.0) 0.0 (1.0) 0.06 (0.0) 0.07 (0.0) 0.02 (0.01)
3 0.03 (0.0) 0.02 (0.05) 0.06 (0.0) 0.07 (0.0) 0.02 (0.02)
4 0.02 (0.0) 0.02 (0.01) 0.08 (0.0) 0.08 (0.0) 0.03 (0.0)
5 0.01 (0.38)* 0.01 (0.47) 0.06 (0.0) 0.08 (0.0) 0.03 (0.0)
6 -0.01 (0.11)* -0.0 (0.68)* 0.05 (0.0)* 0.07 (0.0) 0.02 (0.01)
7 0.01 (0.07)* 0.0 (0.96) 0.05 (0.0)* 0.08 (0.0) 0.02 (0.01)
8 0.0 (0.85)* -0.01 (0.33)* 0.04 (0.0)* 0.07 (0.0) 0.02 (0.04)
9 0.02 (0.01) -0.01 (0.18)* 0.03 (0.0)* 0.05 (0.0)* -0.0 (0.77)*
10 0.01 (0.48)* -0.02 (0.02)* 0.04 (0.0)* 0.04 (0.0)* -0.02 (0.0)*
11 0.02 (0.0) -0.02 (0.04)* 0.03 (0.0)* 0.02 (0.03)* -0.03 (0.0)*
12 0.01 (0.1)* -0.01 (0.08)* -0.02 (0.01)* 0.02 (0.04)* -0.02 (0.0)*
13 0.01 (0.5)* -0.0 (0.66)* -0.05 (0.0)* 0.03 (0.0)* -0.04 (0.0)*
14 - - 0.05 (0.0)* - -
Table 5: Layer-wise RSA values (Spearman correlations) for the sentence condition and visual network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LLaVA IDEFICS2 Llama3 Mistral
1 -0.05 (0.0)* -0.01 (0.39)* -0.05 (0.0)* -0.01 (0.39)*
2 -0.02 (0.03)* 0.02 (0.01)* -0.03 (0.0)* 0.02 (0.02)*
3 0.0 (0.71) -0.01 (0.23)* -0.02 (0.0)* -0.01 (0.48)*
4 0.01 (0.5) -0.03 (0.0)* 0.0 (0.99) -0.05 (0.0)*
5 -0.0 (0.83) -0.03 (0.0)* -0.0 (0.91) -0.04 (0.0)*
6 -0.03 (0.0)* 0.02 (0.05)* -0.02 (0.03) 0.01 (0.45)*
7 -0.05 (0.0)* -0.01 (0.29)* -0.04 (0.0)* -0.02 (0.02)*
8 -0.05 (0.0)* -0.03 (0.0)* -0.04 (0.0)* -0.02 (0.0)*
9 -0.07 (0.0)* -0.06 (0.0)* -0.06 (0.0)* -0.04 (0.0)*
10 -0.07 (0.0)* -0.07 (0.0)* -0.07 (0.0)* -0.05 (0.0)*
11 -0.08 (0.0)* -0.06 (0.0)* -0.08 (0.0)* -0.05 (0.0)*
12 -0.07 (0.0)* -0.05 (0.0)* -0.08 (0.0)* -0.05 (0.0)*
13 -0.08 (0.0)* -0.05 (0.0)* -0.09 (0.0)* -0.05 (0.0)*
14 -0.09 (0.0)* -0.06 (0.0)* -0.09 (0.0)* -0.06 (0.0)*
15 -0.09 (0.0)* -0.04 (0.0)* -0.09 (0.0)* -0.04 (0.0)*
16 -0.06 (0.0)* -0.04 (0.0)* -0.04 (0.0)* -0.04 (0.0)*
17 -0.05 (0.0)* -0.02 (0.01)* -0.01 (0.16) -0.02 (0.01)*
18 -0.07 (0.0)* -0.04 (0.0)* -0.02 (0.01)* -0.03 (0.0)*
19 -0.04 (0.0)* -0.03 (0.0)* 0.0 (0.81) -0.03 (0.0)*
20 -0.04 (0.0)* 0.0 (0.69)* 0.01 (0.41) 0.02 (0.04)*
21 -0.05 (0.0)* 0.02 (0.0) 0.0 (0.54) 0.04 (0.0)
22 -0.04 (0.0)* 0.03 (0.0) -0.0 (0.63) 0.04 (0.0)
23 -0.02 (0.02)* 0.04 (0.0) 0.0 (0.6) 0.04 (0.0)
24 -0.02 (0.0)* 0.05 (0.0) -0.01 (0.18) 0.04 (0.0)
25 -0.02 (0.01)* 0.04 (0.0) -0.02 (0.02) 0.03 (0.0)
26 -0.02 (0.0)* 0.03 (0.0) -0.03 (0.0)* 0.02 (0.01)
27 -0.03 (0.0)* 0.03 (0.0) -0.03 (0.0)* 0.02 (0.03)*
28 -0.03 (0.0)* 0.02 (0.0) -0.03 (0.0)* 0.01 (0.18)*
29 -0.03 (0.0)* 0.01 (0.26)* -0.04 (0.0)* -0.0 (0.81)*
30 -0.02 (0.0)* -0.01 (0.33)* -0.03 (0.0)* -0.01 (0.17)*
31 -0.03 (0.0)* -0.02 (0.05)* -0.04 (0.0)* -0.01 (0.36)*
32 -0.02 (0.04)* -0.02 (0.05)* -0.04 (0.0)* -0.01 (0.11)*
33 0.01 (0.1) -0.01 (0.52)* 0.0 (0.57) -0.01 (0.52)*
Table 6: Layer-wise RSA values (Spearman correlations) for the sentence condition and LH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LLaVA IDEFICS2 Llama3 Mistral
1 0.01 (0.2)* 0.01 (0.17)* 0.01 (0.26)* 0.01 (0.17)*
2 -0.01 (0.26)* -0.05 (0.0)* 0.0 (0.86)* -0.01 (0.07)*
3 0.03 (0.0)* 0.01 (0.12)* 0.02 (0.05)* 0.01 (0.14)*
4 0.06 (0.0) -0.0 (1.0)* 0.04 (0.0) -0.0 (0.83)*
5 0.03 (0.0) -0.0 (0.6)* 0.03 (0.0) -0.02 (0.04)*
6 0.04 (0.0) 0.01 (0.08)* 0.04 (0.0) 0.01 (0.18)*
7 0.01 (0.12)* 0.0 (0.68)* 0.01 (0.06)* -0.0 (0.93)*
8 0.01 (0.47)* -0.01 (0.12)* -0.0 (0.74)* -0.01 (0.08)*
9 -0.0 (0.71)* -0.01 (0.11)* -0.01 (0.16)* -0.01 (0.5)*
10 -0.0 (0.88)* -0.03 (0.0)* -0.02 (0.04)* -0.03 (0.0)*
11 0.0 (0.66)* -0.03 (0.0)* -0.01 (0.16)* -0.02 (0.0)*
12 -0.0 (0.77)* -0.03 (0.0)* -0.02 (0.02)* -0.02 (0.01)*
13 -0.01 (0.12)* -0.02 (0.0)* -0.02 (0.01)* -0.02 (0.0)*
14 -0.02 (0.02)* -0.02 (0.01)* -0.02 (0.01)* -0.02 (0.03)*
15 -0.04 (0.0)* -0.02 (0.03)* -0.04 (0.0)* -0.02 (0.02)*
16 0.01 (0.09)* -0.03 (0.0)* 0.02 (0.03)* -0.03 (0.0)*
17 0.01 (0.28)* -0.02 (0.04)* 0.01 (0.37)* -0.02 (0.01)*
18 -0.0 (0.9)* -0.01 (0.11)* -0.02 (0.05)* -0.01 (0.07)*
19 0.01 (0.49)* -0.04 (0.0)* -0.02 (0.02)* -0.04 (0.0)*
20 0.02 (0.0)* 0.01 (0.25)* -0.01 (0.13)* 0.0 (0.61)*
21 0.02 (0.0)* 0.01 (0.06)* -0.01 (0.49)* 0.01 (0.14)*
22 0.03 (0.0)* 0.03 (0.0)* 0.0 (0.67)* 0.02 (0.0)*
23 0.05 (0.0) 0.03 (0.0)* 0.02 (0.02) 0.03 (0.0)*
24 0.05 (0.0) 0.04 (0.0)* 0.02 (0.04)* 0.03 (0.0)*
25 0.04 (0.0) 0.04 (0.0)* 0.01 (0.29)* 0.03 (0.0)*
26 0.04 (0.0) 0.04 (0.0)* 0.0 (0.68)* 0.03 (0.0)*
27 0.04 (0.0) 0.05 (0.0)* 0.0 (0.57)* 0.04 (0.0)*
28 0.04 (0.0) 0.05 (0.0)* 0.0 (0.76)* 0.04 (0.0)*
29 0.04 (0.0) 0.04 (0.0)* -0.0 (0.58)* 0.03 (0.0)*
30 0.04 (0.0) 0.05 (0.0) -0.0 (0.99)* 0.04 (0.0)*
31 0.04 (0.0) 0.06 (0.0) -0.0 (0.54)* 0.05 (0.0)*
32 0.04 (0.0) 0.06 (0.0) -0.0 (0.62)* 0.06 (0.0)
33 0.03 (0.0)* 0.08 (0.0) 0.01 (0.35)* 0.08 (0.0)
Table 7: Layer-wise RSA values (Spearman correlations) for the sentence condition and RH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LLaVA IDEFICS2 Llama3 Mistral
1 -0.01 (0.08)* 0.01 (0.27)* -0.02 (0.05)* 0.01 (0.27)*
2 -0.02 (0.04)* -0.01 (0.2)* -0.02 (0.01)* 0.0 (0.75)*
3 0.01 (0.07)* 0.02 (0.0)* -0.01 (0.47)* 0.02 (0.01)*
4 0.03 (0.0)* -0.01 (0.28)* 0.0 (0.95)* -0.02 (0.06)*
5 0.01 (0.11)* -0.01 (0.08)* 0.0 (0.71)* -0.02 (0.0)*
6 0.01 (0.07)* 0.0 (0.93)* 0.02 (0.02) -0.0 (0.86)*
7 -0.0 (0.54)* -0.0 (0.6)* -0.0 (0.78)* -0.01 (0.32)*
8 -0.01 (0.08)* -0.02 (0.05)* -0.02 (0.04)* -0.01 (0.17)*
9 -0.02 (0.0)* -0.03 (0.0)* -0.03 (0.0)* -0.01 (0.1)*
10 -0.03 (0.0)* -0.05 (0.0)* -0.04 (0.0)* -0.03 (0.0)*
11 -0.03 (0.0)* -0.05 (0.0)* -0.05 (0.0)* -0.03 (0.0)*
12 -0.03 (0.0)* -0.03 (0.0)* -0.05 (0.0)* -0.03 (0.0)*
13 -0.03 (0.0)* -0.03 (0.0)* -0.05 (0.0)* -0.03 (0.0)*
14 -0.06 (0.0)* -0.03 (0.0)* -0.06 (0.0)* -0.03 (0.0)*
15 -0.06 (0.0)* -0.03 (0.0)* -0.06 (0.0)* -0.03 (0.0)*
16 -0.02 (0.01)* -0.04 (0.0)* -0.01 (0.08)* -0.04 (0.0)*
17 -0.01 (0.15)* -0.02 (0.01)* -0.0 (0.81)* -0.02 (0.0)*
18 -0.03 (0.0)* -0.03 (0.0)* -0.02 (0.05)* -0.02 (0.0)*
19 0.0 (0.54)* -0.03 (0.0)* 0.01 (0.29)* -0.03 (0.0)*
20 0.03 (0.0)* 0.03 (0.0)* 0.02 (0.01) 0.03 (0.0)*
21 0.02 (0.01)* 0.04 (0.0)* 0.02 (0.0) 0.04 (0.0)*
22 0.03 (0.0) 0.06 (0.0) 0.03 (0.0) 0.06 (0.0)
23 0.05 (0.0) 0.08 (0.0) 0.04 (0.0) 0.07 (0.0)
24 0.06 (0.0) 0.09 (0.0) 0.04 (0.0) 0.07 (0.0)
25 0.06 (0.0) 0.09 (0.0) 0.03 (0.0) 0.07 (0.0)
26 0.05 (0.0) 0.09 (0.0) 0.02 (0.0) 0.07 (0.0)
27 0.05 (0.0) 0.09 (0.0) 0.02 (0.01) 0.07 (0.0)
28 0.05 (0.0) 0.08 (0.0) 0.02 (0.05)* 0.06 (0.0)
29 0.04 (0.0) 0.06 (0.0) 0.01 (0.15)* 0.05 (0.0)*
30 0.05 (0.0) 0.06 (0.0)* 0.01 (0.12)* 0.05 (0.0)
31 0.03 (0.0)* 0.05 (0.0)* -0.01 (0.36)* 0.05 (0.0)
32 0.02 (0.0)* 0.05 (0.0)* -0.02 (0.06)* 0.04 (0.0)*
33 0.01 (0.31)* 0.06 (0.0)* 0.0 (0.9)* 0.06 (0.0)
Table 8: Layer-wise RSA values (Spearman correlations) for the sentence condition and visual network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LXMERT VisualBERT BERT LLaVA IDEFICS2 Llama3 Mistral
1 0.23 (0.0)* 0.33 (0.0) -0.01 (0.11)* 0.04 (0.0)* -0.05 (0.0)* -0.11 (0.0)* -0.12 (0.0)*
2 0.23 (0.0)* 0.31 (0.0) 0.11 (0.0) 0.05 (0.0)* 0.06 (0.0)* -0.1 (0.0)* -0.1 (0.0)*
3 0.24 (0.0)* 0.32 (0.0) 0.06 (0.0)* -0.08 (0.0)* -0.0 (0.56)* -0.13 (0.0)* -0.13 (0.0)*
4 0.23 (0.0)* 0.32 (0.0) 0.07 (0.0)* -0.0 (0.83)* -0.04 (0.0)* -0.11 (0.0)* -0.1 (0.0)*
5 0.23 (0.0)* 0.32 (0.0) 0.01 (0.5)* 0.03 (0.0)* -0.04 (0.0)* -0.15 (0.0)* -0.09 (0.0)*
6 0.26 (0.0)* 0.31 (0.0) -0.07 (0.0)* 0.07 (0.0)* -0.03 (0.0)* -0.13 (0.0)* -0.05 (0.0)*
7 0.26 (0.0)* 0.3 (0.0)* -0.08 (0.0)* 0.1 (0.0)* -0.06 (0.0)* -0.13 (0.0)* -0.06 (0.0)*
8 0.28 (0.0)* 0.31 (0.0) -0.07 (0.0)* 0.09 (0.0)* -0.04 (0.0)* -0.13 (0.0)* -0.04 (0.0)*
9 0.27 (0.0)* 0.31 (0.0)* -0.04 (0.0)* 0.08 (0.0)* -0.06 (0.0)* -0.13 (0.0)* -0.05 (0.0)*
10 0.27 (0.0)* 0.31 (0.0)* -0.03 (0.0)* 0.11 (0.0)* -0.01 (0.08)* -0.13 (0.0)* -0.05 (0.0)*
11 0.27 (0.0)* 0.32 (0.0) -0.02 (0.01)* 0.09 (0.0)* -0.01 (0.06)* -0.13 (0.0)* -0.07 (0.0)*
12 0.28 (0.0)* 0.33 (0.0) 0.02 (0.05)* 0.07 (0.0)* -0.03 (0.0)* -0.13 (0.0)* -0.08 (0.0)*
13 0.27 (0.0)* 0.33 (0.0) 0.01 (0.16)* 0.04 (0.0)* -0.04 (0.0)* -0.13 (0.0)* -0.09 (0.0)*
14 0.33 (0.0) - - 0.07 (0.0)* -0.04 (0.0)* -0.13 (0.0)* -0.1 (0.0)*
15 - - - 0.07 (0.0)* -0.04 (0.0)* -0.13 (0.0)* -0.1 (0.0)*
16 - - - -0.02 (0.03)* -0.1 (0.0)* -0.13 (0.0)* -0.14 (0.0)*
17 - - - -0.02 (0.06)* -0.08 (0.0)* -0.14 (0.0)* -0.16 (0.0)*
18 - - - 0.02 (0.04)* -0.14 (0.0)* -0.14 (0.0)* -0.19 (0.0)*
19 - - - -0.02 (0.01)* -0.1 (0.0)* -0.13 (0.0)* -0.17 (0.0)*
20 - - - -0.03 (0.0)* -0.06 (0.0)* -0.12 (0.0)* -0.15 (0.0)*
21 - - - -0.05 (0.0)* -0.01 (0.08)* -0.11 (0.0)* -0.16 (0.0)*
22 - - - -0.04 (0.0)* 0.04 (0.0)* -0.09 (0.0)* -0.15 (0.0)*
23 - - - 0.03 (0.0)* 0.12 (0.0)* -0.07 (0.0) -0.11 (0.0)*
24 - - - 0.06 (0.0)* 0.14 (0.0) -0.07 (0.0) -0.09 (0.0)*
25 - - - 0.09 (0.0)* 0.15 (0.0) -0.07 (0.0) -0.06 (0.0)*
26 - - - 0.1 (0.0)* 0.16 (0.0) -0.07 (0.0) -0.04 (0.0)*
27 - - - 0.09 (0.0)* 0.16 (0.0) -0.06 (0.0) -0.02 (0.0)*
28 - - - 0.09 (0.0)* 0.15 (0.0) -0.07 (0.0) -0.01 (0.07)*
29 - - - 0.1 (0.0)* 0.14 (0.0)* -0.06 (0.0) -0.01 (0.07)*
30 - - - 0.12 (0.0)* 0.14 (0.0)* -0.07 (0.0) -0.04 (0.0)*
31 - - - 0.18 (0.0) 0.14 (0.0) -0.07 (0.0) -0.04 (0.0)*
32 - - - 0.17 (0.0) 0.14 (0.0)* -0.08 (0.0)* -0.06 (0.0)*
33 - - - 0.1 (0.0)* 0.1 (0.0)* -0.05 (0.0) 0.04 (0.0)
Table 9: Layer-wise RSA values (Spearman correlations) for the picture condition and LH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LXMERT VisualBERT BERT LLaVA IDEFICS2 Llama3 Mistral
1 0.18 (0.0)* 0.37 (0.0) 0.04 (0.0)* 0.03 (0.0)* 0.01 (0.06)* -0.07 (0.0)* -0.06 (0.0)*
2 0.18 (0.0)* 0.36 (0.0) 0.09 (0.0) 0.05 (0.0)* 0.09 (0.0)* -0.05 (0.0)* -0.04 (0.0)*
3 0.18 (0.0)* 0.36 (0.0) 0.05 (0.0)* -0.01 (0.1)* 0.04 (0.0)* -0.07 (0.0)* -0.08 (0.0)*
4 0.17 (0.0)* 0.36 (0.0) 0.06 (0.0)* 0.04 (0.0)* 0.01 (0.06)* -0.06 (0.0)* -0.06 (0.0)*
5 0.16 (0.0)* 0.35 (0.0) 0.01 (0.38)* 0.05 (0.0)* 0.02 (0.01)* -0.08 (0.0)* -0.04 (0.0)*
6 0.19 (0.0)* 0.35 (0.0)* -0.04 (0.0)* 0.08 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.02 (0.0)*
7 0.19 (0.0)* 0.35 (0.0)* -0.03 (0.0)* 0.12 (0.0)* 0.03 (0.0)* -0.07 (0.0)* -0.04 (0.0)*
8 0.19 (0.0)* 0.35 (0.0) -0.02 (0.03)* 0.12 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.02 (0.02)*
9 0.18 (0.0)* 0.35 (0.0) 0.01 (0.12)* 0.12 (0.0)* 0.03 (0.0)* -0.07 (0.0)* -0.02 (0.01)*
10 0.2 (0.0)* 0.34 (0.0)* 0.01 (0.06)* 0.15 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.02 (0.05)*
11 0.22 (0.0)* 0.34 (0.0)* 0.02 (0.03)* 0.13 (0.0)* 0.05 (0.0)* -0.07 (0.0)* -0.02 (0.0)*
12 0.23 (0.0)* 0.34 (0.0)* 0.04 (0.0)* 0.12 (0.0)* 0.03 (0.0)* -0.07 (0.0)* -0.04 (0.0)*
13 0.23 (0.0)* 0.33 (0.0)* 0.04 (0.0)* 0.09 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.04 (0.0)*
14 0.26 (0.0) - - 0.1 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.04 (0.0)*
15 - - - 0.09 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.04 (0.0)*
16 - - - 0.04 (0.0)* 0.01 (0.36)* -0.07 (0.0)* -0.06 (0.0)*
17 - - - 0.03 (0.0)* 0.01 (0.32)* -0.09 (0.0)* -0.07 (0.0)*
18 - - - 0.06 (0.0)* -0.03 (0.0)* -0.07 (0.0)* -0.08 (0.0)*
19 - - - 0.05 (0.0)* -0.04 (0.0)* -0.07 (0.0)* -0.1 (0.0)*
20 - - - 0.04 (0.0)* 0.0 (0.78)* -0.07 (0.0)* -0.06 (0.0)*
21 - - - 0.04 (0.0)* 0.04 (0.0)* -0.07 (0.0)* -0.05 (0.0)*
22 - - - 0.05 (0.0)* 0.06 (0.0)* -0.05 (0.0)* -0.07 (0.0)*
23 - - - 0.07 (0.0)* 0.1 (0.0) -0.05 (0.0)* -0.05 (0.0)*
24 - - - 0.08 (0.0)* 0.11 (0.0) -0.04 (0.0)* -0.03 (0.0)*
25 - - - 0.09 (0.0)* 0.12 (0.0) -0.04 (0.0)* -0.01 (0.11)*
26 - - - 0.1 (0.0)* 0.12 (0.0) -0.04 (0.0)* 0.0 (0.99)*
27 - - - 0.1 (0.0)* 0.12 (0.0) -0.04 (0.0)* 0.01 (0.29)
28 - - - 0.1 (0.0)* 0.1 (0.0) -0.04 (0.0)* 0.01 (0.09)
29 - - - 0.11 (0.0)* 0.09 (0.0)* -0.04 (0.0)* 0.01 (0.1)
30 - - - 0.12 (0.0)* 0.09 (0.0)* -0.04 (0.0)* -0.0 (0.79)*
31 - - - 0.15 (0.0)* 0.11 (0.0) -0.04 (0.0)* -0.0 (0.8)*
32 - - - 0.16 (0.0) 0.1 (0.0) -0.04 (0.0)* -0.01 (0.2)*
33 - - - 0.18 (0.0) 0.09 (0.0)* 0.03 (0.0) 0.03 (0.0)
Table 10: Layer-wise RSA values (Spearman correlations) for the picture condition and RH language network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
Layer LXMERT VisualBERT BERT LLaVA IDEFICS2 Llama3 Mistral
1 0.2 (0.0)* 0.48 (0.0) -0.01 (0.36)* -0.02 (0.02)* 0.01 (0.08)* -0.07 (0.0)* -0.04 (0.0)*
2 0.2 (0.0)* 0.47 (0.0) 0.08 (0.0) 0.04 (0.0)* 0.1 (0.0)* -0.03 (0.0)* -0.02 (0.0)*
3 0.21 (0.0)* 0.46 (0.0)* 0.06 (0.0) -0.0 (0.72)* 0.05 (0.0)* -0.04 (0.0)* -0.09 (0.0)*
4 0.2 (0.0)* 0.46 (0.0)* 0.06 (0.0) 0.04 (0.0)* 0.05 (0.0)* -0.03 (0.0)* -0.05 (0.0)*
5 0.19 (0.0)* 0.45 (0.0)* 0.02 (0.02)* 0.06 (0.0)* 0.04 (0.0)* -0.05 (0.0)* -0.04 (0.0)*
6 0.21 (0.0)* 0.43 (0.0)* -0.03 (0.0)* 0.08 (0.0)* 0.05 (0.0)* -0.05 (0.0)* 0.0 (0.72)*
7 0.22 (0.0)* 0.42 (0.0)* -0.04 (0.0)* 0.11 (0.0)* 0.03 (0.0)* -0.05 (0.0)* -0.02 (0.03)*
8 0.22 (0.0)* 0.43 (0.0)* -0.02 (0.01)* 0.11 (0.0)* 0.02 (0.04)* -0.05 (0.0)* 0.01 (0.35)*
9 0.22 (0.0)* 0.42 (0.0)* 0.02 (0.03)* 0.11 (0.0)* 0.01 (0.07)* -0.05 (0.0)* -0.0 (0.62)*
10 0.22 (0.0)* 0.41 (0.0)* -0.0 (0.91)* 0.14 (0.0)* 0.03 (0.0)* -0.05 (0.0)* -0.01 (0.46)*
11 0.25 (0.0)* 0.4 (0.0)* -0.01 (0.48)* 0.13 (0.0)* 0.03 (0.0)* -0.05 (0.0)* -0.02 (0.01)*
12 0.27 (0.0) 0.39 (0.0)* 0.02 (0.0)* 0.12 (0.0)* 0.02 (0.01)* -0.05 (0.0)* -0.02 (0.0)*
13 0.27 (0.0) 0.36 (0.0)* 0.03 (0.0)* 0.08 (0.0)* 0.03 (0.0)* -0.05 (0.0)* -0.03 (0.0)*
14 0.29 (0.0) - - 0.1 (0.0)* 0.03 (0.0)* -0.05 (0.0)* -0.04 (0.0)*
15 - - - 0.08 (0.0)* 0.02 (0.0)* -0.05 (0.0)* -0.04 (0.0)*
16 - - - 0.04 (0.0)* -0.01 (0.43)* -0.05 (0.0)* -0.06 (0.0)*
17 - - - 0.05 (0.0)* 0.0 (0.84)* -0.05 (0.0)* -0.06 (0.0)*
18 - - - 0.04 (0.0)* -0.03 (0.0)* -0.05 (0.0)* -0.09 (0.0)*
19 - - - 0.05 (0.0)* -0.01 (0.12)* -0.04 (0.0)* -0.08 (0.0)*
20 - - - 0.02 (0.01)* 0.0 (0.96)* -0.05 (0.0)* -0.08 (0.0)*
21 - - - 0.02 (0.01)* 0.07 (0.0)* -0.05 (0.0)* -0.05 (0.0)*
22 - - - 0.04 (0.0)* 0.08 (0.0)* -0.02 (0.01)* -0.08 (0.0)*
23 - - - 0.07 (0.0)* 0.11 (0.0) -0.0 (0.6)* -0.06 (0.0)*
24 - - - 0.09 (0.0)* 0.12 (0.0) -0.0 (0.98)* -0.05 (0.0)*
25 - - - 0.1 (0.0)* 0.13 (0.0) 0.0 (0.63)* -0.03 (0.0)*
26 - - - 0.1 (0.0)* 0.13 (0.0) 0.0 (0.57)* -0.02 (0.03)*
27 - - - 0.11 (0.0)* 0.13 (0.0) 0.01 (0.29)* -0.0 (0.67)*
28 - - - 0.11 (0.0)* 0.11 (0.0) 0.01 (0.41)* 0.01 (0.4)*
29 - - - 0.12 (0.0)* 0.1 (0.0)* 0.01 (0.41)* 0.01 (0.19)*
30 - - - 0.13 (0.0)* 0.09 (0.0)* 0.0 (0.59)* 0.0 (0.97)*
31 - - - 0.17 (0.0) 0.11 (0.0)* 0.0 (0.71)* -0.01 (0.4)*
32 - - - 0.17 (0.0) 0.1 (0.0)* -0.0 (0.67)* -0.02 (0.05)*
33 - - - 0.17 (0.0) 0.08 (0.0)* 0.08 (0.0) 0.05 (0.0)
Table 11: Layer-wise RSA values (Spearman correlations) for the picture condition and visual network, with $p$-values indicating the significance of the correlation (i.e., whether it is different from 0) in brackets. Note that asterisks indicate whether each correlation is statistically significantly different from the correlation achieved by the best layer (boldfaced value).
[…]
Picture Condition Ablation: Additional Comparison
To facilitate comparison between VLMs’ brain alignment in the picture condition when the image is passed vs. when it is not, we visualise this information in Figure 7.
Figure 7: Initial VLM results from RSA analysis in the picture condition vs. results obtained when passing only concept words (without images).
Acknowledgements.
Anna Bavaresco and Raquel Fernández are funded by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (ERC Consolidator Grant 819455). Marianne de Heer Kloots is funded by the Netherlands Organization for Scientific Research (NWO), through a Gravitation Grant 024.001.006 to the Language in Interaction Consortium.